OpenStudy (anonymous):

(3/10y^2)(1/3y^2)^2 simplify

OpenStudy (anonymous):

3/(10y^2.9y^4) =3/90y^6 =3/90(1/y^6) 3/90(y^-6)

OpenStudy (anonymous):

\[(3/10y^{2})(1/3y^{2})^{2}\]\[(3/10y^{2})(1/9y^{4})\]\[1/30y^{6}\]

jhonyy9 (jhonyy9):

\[(3/10y^{2})(1/3y ^{2})^{2}=(3/10y ^{2})(1/9y ^{4})=3/90y ^{6}=1/30y ^{6}\]